### v2

@@ 1 Introduction
Experiments with hydrogen atoms were at the core of many discoveries in quantum physics and precision spectroscopy in the history of physics. The dispersion of atomic velocities due to the thermal motion is always a limitation for reaching highest resolution and accuracy. Obtaining an intense beam of slow hydrogen atoms or cooling trapped hydrogen gas is important for further improvements in optical, hyperfine and gravitational spectroscopy [1] and for potential discoveries of new phenomena. Laser cooling routinely used for many other atoms and molecules is very difficult for hydrogen due its low mass and VUV range of the required excitation from the ground electronic state. Magnetic trapping and evaporative cooling are the methods which were successfully used for reaching quantum degeneracy and BEC in H gas in the end of the last century [2]. Revisiting this techniques, we have recently built a large magnetic trap which is planned to be used for a variety of experiments with ultra-cold hydrogen atoms by the GRASIAN collaboration [3]. Loading such a trap requires an intense source of low-field seeking (lfs) atoms at sub-Kelvin temperatures. Here, we describe the design and operation of such a source and demonstrate its efficiency for a loading the magnetic trap. The source is based on a cryogenic dissociator operating at 0.65 K and a hydrogen transfer line covered by superfluid helium film and thermally anchored to a different stages of a dilution refrigerator. A gradient of magnetic field separates low-field and high-field seeking (hfs) atoms. The source can be used as a stand-alone beam of slow hydrogen atoms. Elastic scattering from a superfluid helium covered surfaces can be utilized for further manipulation of the beam velocity and intensity.
In Section 2 of this paper, we present a brief review of the cold atomic hydrogen sources and techniques used for various experiments, in Section 3 we describe shortly our experimental setup with a magnetic trap and a dilution refrigerator, Section 4 contains a description of the low-temperature hydrogen dissociator and hydrogen transfer line with the thermal accommodators. We present results of the tests of the source operation and loading the magnetic trap as well as of a decay of the trapped gas in the Section 5. Finally, we analyze further improvements of the source for getting colder atoms with higher flux and suggest some ideas on the manipulation of the cold H beam using quantum reflection from a superfluid helium covered surface.
@@ 2 Background
In a pioneering work of I. Silvera and J. Walraven [4] an intense cryogenic beam of H was obtained by dissociating H2 at room temperature in the RF discharge and further thermal accommodation in a nozzle cooled by liquid helium. The atoms were cooled to K in collisions with the nozzle walls and a fairly large flux of atoms/s was demonstrated. Such a cold nozzle technique was then used for high precision optical spectroscopy with atomic hydrogen [5, 6, 7], microwave spectroscopy [8] and experiments on quantum reflection of slow hydrogen beam [9, 10]. In these experiments the typical thermal distribution of atomic velocities after the nozzle corresponded to the temperature of K with a velocity peak at m/s. A special time-of-flight technique allowed a selection of slowest atoms in this distribution down to 50-60 m/s [11, 10]. Further reduction was not possible due to the low intensity of the atomic flux and insufficient sensitivity of the detection system.
@@ 4.3 Hydrogen transfer line
The magnetic field configuration is rather complex between the dissociator coil and the octupole magnets since the field decreases not only in magnitude but also changes the direction. The actual field profiles along the transfer line were evaluated using the BiotSavart program [36] and used in the design of each section. The length of each section was done as short as possible, however taking into account that H atoms should experience on the average 1-2 sticking collisions with the wall. To increase the number of collisions, the 90∘ turns were included. This also prevents a direct access of ”hot” atoms into the trapping cell. The turns as well help to fit all sections in a fairly small space available inside the dissociator coil and octupole magnets. In Fig. 3 we present a schematic plot of the magnetic field and temperatures changes across different section of the transfer line. One can see that the magnetic field drops by T between the dissociator and Ac2 and the temperature decreases from 600 to 320 mK on a short length of cm. This provides effective separation of the lfs and hfs before the lfs enter the CI, the coldest and longest part of the transfer line.
The last part of the transfer line (CI) is made of a 90∘ turn at the top (see Fig. 2 followed by a 3 mm i.d. copper tube of the length cm. The lower end of the tube is covered by a short end cap spraying incoming atoms with the nearly 45∘ with respect to the trap axis. The end of the CI extends for about 5 cm below the upper flange of the trapping cell, in the region with magnetic field of T. This way we can load H gas into the trap at the level of roughly half of the trapping barrier and reduce the kinetic energy gained by the atoms when they settle in the trap. The incoming gas temperature of 130 mK is comparable with the magnetic energy difference of 0.2 K between the end of the CI and the trap bottom and a factor of smaller than the 540 mK trap barrier. This provides effective trapping of a large fraction of the incoming gas. The average number of sticking collisions in the CI is , which is somewhat larger than needed. There is a danger of enhanced recombination on the walls of the CI, since the surface residence time increases exponentially with decreasing of temperature [37]. However, in this part of the line the gas is well polarized. The magnetic compression factor separating the lfs from the hfs is exp. We installed a heater on the CI outside the trapping cell, which can be used to warm up this section in order to decrease recombination.
@@ 4.4 Sample cell for the trapped gas
The hydrogen sample cell (SC) is a cylinder made of G10 Glass Epoxy Laminate tube [39] with inner diameter of 105 mm and wall thickness of 1.6 mm (Fig. 4). It fits inside the IPT with a mm gap to the inner surface of the octupole system. Centering and control of the gap between SC and IPT is performed with 8 PEEK rods screwed through the cores of the race track coils of the octupole magnets and pressing onto the SC tube with sharpened ends. Cooling of the SC is provided by the thermal links made of ribbons of 0.15 mm diameter insulated copper wire glued to the outer surface of the G10 tube. This helps reducing the heating of the SC by eddy currents during ramps of magnetic field. SC top and bottom end plates are manufactured from a 6 mm thick G10 sheet. Copper wire ribbons are glued to the copper plates at the SC top and bottom. Separate copper bars connect these plates to the mixing chamber of the DR. The bar for the SC bottom is placed outside the IPT. Temperature sensors and heaters are fixed to the top and bottom SC plates. This allows independent control of the SC top and bottom temperatures.
@@
For reaching better sensitivity, we installed bolometer at the bottom of the SC (Fig. 4). The bolometer has two temperature sensors located on different sides of a dielectric substrate. The sensors are carbon resistors made by painting Aquadag (colloidal solution of carbon particles in water [41]) between the thin electrodes sputtered on the substrate. We used a 0.1 mm thick sapphire sheet for the bolometer substrate with an area of cm2. The substrate is thermally coupled to the SC wall by a 50 m thick superconductive wire of cm length. Temperature sensors on both sides of the substrate provide the possibility of an absolute power measurement. For this, one Aquadag resistor is used as a heater using a high excitation current. A second resistor serves as a thermometer, measured with a very low excitation. This also allows stabilization of the substrate temperature at a given value avoiding uncertainties with the overheating effect of the Aquadag layer with respect to the helium film on it.
Bolometers are capable of detecting power changes at a pW level. However, they detect only a fraction of the total recombination heat equal to the ratio of the geometrical area of the bolometer to the area of the inner surface of the SC. This is for our SC. Taking this into account, the bolometers are still a factor of 10-100 more sensitive than the temperature controllers described above. The presence of H gas provides an extra cooling channel via a thermal accommodation of the gas [42]. This effect depends on the gas density. As we shall see below, this makes interpretation of the bolometer signals rather complicated.
@@ 5.1 Loading the sample cell with H gas
In the first loading experiments, we charged the IPT coils to the currents which provide nearly the same trapping barrier for the axial and radial confinement. The currents which we fixed as nominal to operate the IPT safely with minimal risk of a quench were 120, 95, and 32 A for the octupole system, lower pinch, and upper pinch accordingly. We denote this configuration as the full ”100” IPT charge. Later we will describe experiments where the IPT was charged to smaller fields. In all cases, the dissociator coil was charged to a nominal value of 95 A corresponding to T. The full charge in the IPT provides a minimum trapping barrier of T (0.54 K) in the so-called ”leakage points” [3] where the field of the octupole magnets is partially canceled by the field of the pinch coils. In the other regions the field is higher, reaching a maximum of T. Measurements were performed at SC temperatures in the range of 119…243 mK. The mixing chamber was stabilized at the lowest possible temperature which the DR could provide under a certain heat load from the SC, the transfer line components and the dissociator running the RF discharge. Typically the MC operated in the range 40-70 mK. The bolometer at the SC bottom was stabilized to 315 mK.
@@
After stabilization of the SC temperature, we turned on the RF discharge in the dissociator and recorded voltages of the SC temperature controllers and the bolometer stabilization power. The power delivered to the dissociator was varied by changing the power of the RF oscillator by 6 dB or changing the duty cycle of its pulses. The average power absorbed in the dissociator running RF discharge was mW. Normally we used pulses of H/L = 0.8/(20…40) ms with ”H” RF power ON and ”L” RF power OFF times. After 10-15 minutes of loading we stopped the dissociator and monitored the decay of the trapped gas. In Fig. 5 we presented results of such an experiment for the fully energized trap at the SC temperature of 134 and 167 mK. One can see that running the dissociator provides extra recombination power, which leads to a decrease of the temperature controller power required for stabilization of the SC. The effect is saturated after s of accumulation with a power change of 25 W. Taking into account the recombination energy of J per atom, we evaluate that a flux of atoms/s is coming into the SC. For the mK we first run the dissociator pulsing with H/L=0.8/40 ms, then increased the duty cycle by reducing L to 30 ms. This led to an increase of the detected atomic flux by about a factor of 2. Finally, the temperature controllers were not capable to keep stabilizing the SC since the recombination power exceeded the limit set by the temperature difference between the MC and SC (W in this case) with the dissociator OFF. The flux of atoms evaluated at 134 mK is about the same as for 167 mK.
@@
In order to check that the observed change of the SC heating is related with the flux of atoms followed by the recombination, we run a control experiment when there were no molecular hydrogen uploaded into the dissociator. Running the discharge under the same conditions, we observed some heating of the dissociator and accommodators. No change of the heating power in the SC was detected within the noise, at the level of W. In this regime some part of the superfluid helium film inside the dissociator is evaporated and re-condensed in the accommodators before reaching the SC. The large difference in the heating due to the hydrogen flux demonstrates a high efficiency of the transfer line and a proper design of the accommodators.
@@ 5.2 Bolometer data
where A is the bolometer surface area, is the thermal velocity of atoms, is the Boltzmann constant and is the thermal accommodation coefficient [38]. Using this equation and the ”overshooting” power of the bolometer nW we evaluate that the density of the H gas near the bolometer is cm-3. The magnetic field in the location of the bolometer is T and the magnetic compression factor for the gas located in the trap center is for a temperature of 243 mK. As we shall see below, the maximum density of lfs atoms in the trap center at 243 mK is about cm-3. Then, near the wall the density of lfs should be cm-3, much smaller than estimated from the bolometer cooling. Therefore, the gas which provides cooling cannot be in the lfs state, but is most likely in the hfs. At so high temperature of the SC this conclusion does not look surprising. The magnetic separation cannot effectively separate electron spin states and the recombination at the wall is slow enough to allow substantial density of H gas. The latter is also seen from very slow decay of the bolometer signal after switching off the dissociator.
@@ 5.5 Loading at various trap depths and SC temperatures
We performed a series of measurements with the IPT coils energized to 100, 75, 50 and for the configuration ”0”. In the latter case the trap is open at the bottom while there is T field in its upper part. The atoms are not confined magnetically at the lower part of the SC and freely interact with the SC walls. In Fig. 9 we present results of the temperature controller signal during accumulation of the H gas for the above mentioned magnetic filed configurations in the IPT. At each configuration we performed a series of measurements for various SC temperatures in the range mK. In Fig. 10 we present results of the accumulation for some values of temperature in this range and fully charged IPT.

### v3

@@ 1 Introduction
Experiments with hydrogen atoms were at the core of many discoveries in quantum physics and precision spectroscopy in the history of physics. The dispersion of atomic velocities due to the thermal motion is always a limitation for reaching highest resolution and accuracy. Obtaining an intense beam of slow hydrogen atoms or cooling trapped hydrogen gas is important for further improvements in optical, hyperfine and gravitational spectroscopy [1] and for potential discoveries of new phenomena. Laser cooling routinely used for many other atoms and molecules is very difficult for hydrogen due its low mass and VUV range of the required excitation from the ground electronic state. Magnetic trapping and evaporative cooling are the methods which were successfully used for reaching quantum degeneracy and BEC in H gas in the end of the last century [2]. Revisiting this techniques, we have recently built a large magnetic trap which is planned to be used for a variety of experiments with ultra-cold hydrogen atoms by the GRASIAN collaboration [3]. Loading such a trap requires an intense source of low-field seeking (lfs) atoms at sub-Kelvin temperatures. Here, we describe the design and operation of such a source and demonstrate its efficiency for a loading the magnetic trap. The source is based on a cryogenic dissociator operating at 0.65 K and a hydrogen transfer line covered by superfluid helium film and thermally anchored to different stages of a dilution refrigerator. A gradient of magnetic field separates low-field and high-field seeking (hfs) atoms. The source can be used as a stand-alone beam of slow hydrogen atoms. Elastic scattering from a superfluid helium covered surfaces can be utilized for further manipulation of the beam velocity and intensity.
In Section 2 of this paper, we present a brief review of the cold atomic hydrogen sources and techniques used for various experiments, in Section 3 we describe shortly our experimental setup with a magnetic trap and a dilution refrigerator, Section 4 contains a description of the low-temperature hydrogen dissociator and hydrogen transfer line with the thermal accommodators. We present results of the tests of the source operation and loading the magnetic trap as well as of decay of the trapped gas in the Section 5. Finally, we analyze further improvements of the source for getting colder atoms with higher flux and suggest some ideas on the manipulation of the cold H beam using quantum reflection from a superfluid helium covered surface.
@@ 2 Background
In the pioneering work of I. Silvera and J. Walraven [4] an intense cryogenic beam of H was obtained by dissociating H2 at room temperature in the RF discharge and further thermal accommodation in a nozzle cooled by liquid helium. The atoms were cooled to K in collisions with the nozzle walls and a fairly large flux of atoms/s was demonstrated. Such a cold nozzle technique was then used for high precision optical spectroscopy with atomic hydrogen [5, 6, 7], microwave spectroscopy [8] and experiments on quantum reflection of slow hydrogen beam [9, 10]. In these experiments the typical thermal distribution of atomic velocities after the nozzle corresponded to the temperature of K with a velocity peak at m/s. A special time-of-flight technique allowed a selection of slowest atoms in this distribution down to 50-60 m/s [11, 10]. Further reduction was not possible due to the low intensity of the atomic flux and insufficient sensitivity of the detection system.
@@ 4.3 Hydrogen transfer line
The magnetic field configuration is rather complex between the dissociator coil and the octupole magnets since the field decreases not only in magnitude but also changes the direction. The actual field profiles along the transfer line were evaluated using the BiotSavart program [36] and used in the design of each section. The length of each section was done as short as possible, however taking into account that H atoms should experience on the average 1-2 sticking collisions with the wall. To increase the number of collisions, the 90∘ turns were included. This also prevents a direct access of “hot” atoms into the trapping cell. The turns as well help to fit all sections in a fairly small space available inside the dissociator coil and octupole magnets. In Fig. 3 we present a schematic plot of the magnetic field and temperatures changes across different section of the transfer line. One can see that the magnetic field drops by T between the dissociator and Ac2 and the temperature decreases from 600 to 320 mK on a short length of cm. This provides effective separation of the lfs and hfs before the lfs enter the CI, the coldest and longest part of the transfer line.
The last part of the transfer line (CI) is made of a 90∘ turn at the top (see Fig. 2) followed by a 3 mm i.d. copper tube of the length cm. The lower end of the tube is covered by a short end cap spraying incoming atoms with the nearly 45∘ with respect to the trap axis. The end of the CI extends for about 5 cm below the upper flange of the trapping cell, in the region with magnetic field of T. This way we can load H gas into the trap at the level of roughly half of the trapping barrier and reduce the kinetic energy gained by the atoms when they settle in the trap. The incoming gas temperature of 130 mK is comparable with the magnetic energy difference of 0.2 K between the end of the CI and the trap bottom and a factor of smaller than the 540 mK trap barrier. This provides effective trapping of a large fraction of the incoming gas. The average number of sticking collisions in the CI is , which is somewhat larger than needed. There is a danger of enhanced recombination on the walls of the CI, since the surface residence time increases exponentially with decreasing of temperature [37]. However, in this part of the line the gas is well polarized. The magnetic compression factor separating the lfs from the hfs is exp. We installed a heater on the CI outside the trapping cell, which can be used to warm up this section in order to decrease recombination.
@@ 4.4 Sample cell for the trapped gas
The hydrogen sample cell (SC) is a cylinder made of G10 Glass Epoxy Laminate tube [39] with inner diameter of 105 mm and wall thickness of 1.6 mm (Fig. 4). It fits inside the IPT with a mm gap to the inner surface of the octupole system. Centering and control of the gap between SC and IPT is performed with 8 polyether ether ketone (PEEK) rods screwed through the cores of the race track coils of the octupole magnets and pressing onto the SC tube with sharpened ends. Cooling of the SC is provided by the thermal links made of ribbons of 0.15 mm diameter insulated copper wire glued to the outer surface of the G10 tube. This helps reducing the heating of the SC by eddy currents during ramps of magnetic field. SC top and bottom end plates are manufactured from a 6 mm thick G10 sheet. Copper wire ribbons are glued to the copper plates at the SC top and bottom. Separate copper bars connect these plates to the mixing chamber of the DR. The bar for the SC bottom is placed outside the IPT. Temperature sensors and heaters are fixed to the top and bottom SC plates. This allows independent control of the SC top and bottom temperatures.
@@
For reaching better sensitivity, we installed a bolometer at the bottom of the SC (Fig. 4). The bolometer has two temperature sensors located on different sides of a dielectric substrate. The sensors are carbon resistors made by painting Aquadag (colloidal solution of carbon particles in water [41]) between the thin electrodes sputtered on the substrate. We used a 0.1 mm thick sapphire sheet for the bolometer substrate with an area of cm2. The substrate is thermally coupled to the SC wall by a 50 m thick superconductive wire of cm length. Temperature sensors on both sides of the substrate provide the possibility of an absolute power measurement. For this, one Aquadag resistor is used as a heater using a high excitation current. A second resistor serves as a thermometer, measured with a very low excitation. This also allows stabilization of the substrate temperature at a given value avoiding uncertainties with the overheating effect of the Aquadag layer with respect to the helium film on it.
Bolometers are capable of detecting power changes at a pW level. However, they detect only a fraction of the total recombination heat equal to the ratio of the geometrical area of the bolometer to the area of the inner surface of the SC. This is for our SC. Taking this into account, the bolometers are still a factor of 10-100 more sensitive than the temperature controllers described above. The presence of H gas provides an extra cooling channel via thermal accommodation of the gas [42]. This effect depends on the gas density. As we shall see below, this makes interpretation of the bolometer signals rather complicated.
@@ 5.1 Loading the sample cell with H gas
In the first loading experiments, we charged the IPT coils to the currents which provide nearly the same trapping barrier for the axial and radial confinement. The currents which we fixed as nominal to operate the IPT safely with minimal risk of a quench were 120, 95, and 32 A for the octupole system, lower pinch, and upper pinch accordingly. We denote this configuration as the full “100” IPT charge. Later we will describe experiments where the IPT was charged to smaller fields. In all cases, the dissociator coil was charged to a nominal value of 95 A corresponding to T. The full charge in the IPT provides a minimum trapping barrier of T (0.54 K) in the so-called “leakage points” [3] where the field of the octupole magnets is partially canceled by the field of the pinch coils. In the other regions the field is higher, reaching a maximum of T. Measurements were performed at SC temperatures in the range of 119…243 mK. The mixing chamber was stabilized at the lowest possible temperature which the DR could provide under a certain heat load from the SC, the transfer line components and the dissociator running the RF discharge. Typically the MC operated in the range 40-70 mK. The bolometer at the SC bottom was stabilized to 315 mK.
@@
After stabilization of the SC temperature, we turned on the RF discharge in the dissociator and recorded voltages of the SC temperature controllers and the bolometer stabilization power. The power delivered to the dissociator was varied by changing the power of the RF oscillator by 6 dB or changing the duty cycle of its pulses. The average power absorbed in the dissociator running RF discharge was mW. Normally we used pulses of H/L = 0.8/(20…40) ms with “H” RF power ON and “L” RF power OFF times. After 10-15 minutes of loading we stopped the dissociator and monitored the decay of the trapped gas. In Fig. 5 we presented results of such an experiment for the fully energized trap at the SC temperature of 134 and 167 mK. One can see that running the dissociator provides extra recombination power, which leads to a decrease of the temperature controller power required for stabilization of the SC. The effect is saturated after s of accumulation with a power change of 25 W. Taking into account the recombination energy of J per atom, we evaluate that a flux of atoms/s is coming into the SC. For the mK we first run the dissociator pulsing with H/L=0.8/40 ms, then increased the duty cycle by reducing L to 30 ms. This led to an increase of the detected atomic flux by about a factor of 2. Finally, the temperature controllers were not capable to keep stabilizing the SC since the recombination power exceeded the limit set by the temperature difference between the MC and SC (W in this case) with the dissociator OFF. The flux of atoms evaluated at 134 mK is about the same as for 167 mK.
@@
In order to check that the observed change of the SC heating is related with the flux of atoms followed by the recombination, we run a control experiment when there was no molecular hydrogen uploaded into the dissociator. Running the discharge under the same conditions, we observed some heating of the dissociator and accommodators. No change of the heating power in the SC was detected within the noise, at the level of W. In this regime some part of the superfluid helium film inside the dissociator is evaporated and re-condensed in the accommodators before reaching the SC. The large difference in the heating due to the hydrogen flux demonstrates a high efficiency of the transfer line and a proper design of the accommodators.
@@ 5.2 Bolometer data
where A is the bolometer surface area, is the thermal velocity of atoms, is the Boltzmann constant and is the thermal accommodation coefficient [38]. Using this equation and the “overshooting” power of the bolometer nW we evaluate that the density of the H gas near the bolometer is cm-3. The magnetic field in the location of the bolometer is T and the magnetic compression factor for the gas located in the trap center is for a temperature of 243 mK. As we shall see below, the maximum density of lfs atoms in the trap center at 243 mK is about cm-3. Then, near the wall the density of lfs should be cm-3, much smaller than estimated from the bolometer cooling. Therefore, the gas which provides cooling cannot be in the lfs state, but is most likely in the hfs. At so high temperature of the SC this conclusion does not look surprising. The magnetic separation cannot effectively separate electron spin states and the recombination at the wall is slow enough to allow substantial density of H gas. The latter is also seen from very slow decay of the bolometer signal after switching off the dissociator.
@@ 5.5 Loading at various trap depths and SC temperatures
We performed a series of measurements with the IPT coils energized to 100, 75, 50 and for the configuration “0” when the DS and UP pinch were energized at 100% while all other coils fully discharged (zero field in IPT), so-called open bottom configuration (OB). In the latter case the trap is open at the bottom while there is T field in its upper part. The atoms are not confined magnetically at the lower part of the SC and freely interact with the SC walls. In Fig. 9 we present results of the temperature controller signal during accumulation of the H gas for the above mentioned magnetic field configurations in the IPT. At each configuration we performed a series of measurements for various SC temperatures in the range mK. In Fig. 10 we present results of the accumulation for some values of temperature in this range and fully charged IPT.